# GroUSE: A Benchmark to Evaluate Evaluators in Grounded Question Answering

### v1

@@ Abstract
Sacha Muller  António Loison  Bilel Omrani  Gautier Viaud Illuin Technology { sacha.muller, antonio.loison, bilel.omrani, gautier.viaud }@illuin.tech.fr
@@ 2 Related work
Several prior works have explored methods for evaluating the generator module in RAG systems. Various sets of metrics have been proposed to measure different failure modes, but there is no consensus on a common set of criteria for evaluating grounded question answering. Chen et al. (2024) propose and evaluate 4 abilities required for RAG: noise robustness, negative rejection, information integration and counterfactual robustness. Es et al. (2023) propose two other criteria: faithfulness and answer relevancy. Faithfulness evaluates the factual consistency of the answer given the grounding contexts by decomposing the answer into several statements and calculating the proportion of facts that are supported by the contexts. Answer relevancy measures how well the provided answer addresses the original question. An LLM is used to generate several questions from the answer, the answer relevancy score is then given by the average cosine similarity between dense embeddings of the generated questions and the original question. In Deepeval [^2] https://github.com/confident-ai/deepeval, answer relevancy is computed by using a judge LLM to divide the answer into several atomic facts and computed as the proportion of facts that are relevant to the question. Yu et al. (2024) survey several prior works and propose to add correctness to this pair of metrics, which measures the accuracy of the generated response against a ground truth response. Magesh et al. focus on the legal domain and propose a more fine-grained measure of faithfulness, by distinguishing between the factual accuracy of the response and the validity of the accompanying citations. Thakur et al. (2023) introduce NoMIRACL, a human-labeled dataset of multilingual queries and both relevant and non-relevant subsets to evaluate if the generator correctly refrains from answering with non-relevant passages and correctly recognizes the relevant passage otherwise.
@@ 7. FM7
Table 1: Equivalent failure modes studied and reported in prior works. Existing studies focus on detecting and evaluating a subset of failure modes. For instance FM1 is related to answer relevancy in Es et al. (2023), FM2 is related to negative rejection in Chen et al. (2024) FM6 and FM7 are related to faithfulness and more specifically to correctness and groundedness respectively in Magesh et al. , FM6 is related to citation recall in Gao et al. (2023).
@@ 4.2 Meta-evaluation with unit-testing
We introduce GroUSE (Grounded QA Unitary Scoring of Evaluators), a benchmark consisting of 144 tests divided into 9 sets of 16 tests. All tests within a given set share the same question, with the references and answers slightly modified to fit each of the 16 test types. An additional set of 16 tests is available as a “training set” to assist in engineering the prompt for the judge model being tested. The references are primarily handpicked excerpts from Wikipedia or newspaper articles. The themes of the sets span various domains, including history, science, zoology, cinematography, and the medical field. See Appendix A for details.
@@ 4.4 Enhancing existing frameworks
Each prompt was engineered to fit the specific metric being evaluated, but for all metrics we ask the model to rate two answers, the first one being a reference answer. The model’s reasoning is also guided by the expected format of the JSON output. Following best practices recommendations from Biderman et al. (2024), details about the prompts format and the prompt engineering process are available in Appendix C.
@@
Table 3 shows the performances of various models on GroUSE, across the six metrics, as well as all metrics combined. GPT-4 is the best evaluator, with an overall score of $95\%$, while the best open-source evaluator is Llama-3 70b with a score of $79\%$. The gap between closed and open-source models thus remains wide, with a $15.85\%$ difference on the tests results. Interestingly, Prometheus 2 8x7b does not outperform Mixtral 8x7b, despite Prometheus 2 being specialized in evaluation tasks. It’s worth noting however that Prometheus 2 was trained to predict Likert scores ranging from 1 to 5, whereas our evaluation metrics include boolean and nullable scores, which is outside its intended scope.
@@
Annex B protocol Human annotators 98.61 95.14 98.61 97.92 98.61 99.31 98.03
@@ 5.1 Experimental setup
We finetuned a Llama-3 8b on 1k samples of this dataset, and used the rest as a test set. We refer to Appendix F for details on the finetuning configuration. To measure the model’s progress, we tested its performances on GroUSE. Following Kim et al. (2023), we also report the correlation between GPT-4’s grades and the finetuned model’s grades on the test set. For metrics using a Likert scale, alignment is measured using the Spearman correlation. Other metrics are measured using nullable boolean values, their alignment is evaluated using a three-class macro F1-score.
@@ 5.2 Experimental results
Table 3 presents the pass rate of various judge models on GroUSE, including Llama-3 8b (zero-shot) and our finetuned Llama-3 8b judge model. Finetuning significantly enhances the evaluation capabilities of Llama-3, as evidenced by the substantial improvement in pass rates. Despite extensive prompt engineering and intermediate CoT reasoning (see Appendix C), the non-finetuned Llama-3 8b passes only 40% of the unit tests. However, after finetuning, its pass rate increases to 83%, surpassing all other open-source judges, including Prometheus 2 8x7b (except in the category of faithfulness), despite its smaller size.
@@ Acknowledgments
This research work is supported by Illuin Technology. We would like to express our gratitude to Stanislas Dozias for the insightful discussions and exchange of ideas that contributed to the development of this paper. Special thanks to Quentin Lutz and Manuel Faysse for their valuable feedback and support throughout the research process. Additionally, we would like to thank Noa Grollimund, Florian Muller, Benoît Muller, Max Conti, and Paul Boulenger for their assistance in evaluating the human performance on GroUSE.
@@ Appendix A Unit test characteristics
A detailed listing of the unit tests, including the expected mark for each test, is available in Table 6. The queries of the sets are:
@@
Appendix B Annotation procedures
GroUSE Dataset.
The grounding documents primarily consist of excerpts from Wikipedia, supplemented with manually scraped content from various sources such as news articles, popular science pieces, and medical papers. To simulate retrieval system noise, the references were intentionally altered by truncating sentences, mimicking poorly parsed tables, and including irrelevant headers or footers. To further replicate real-world retrieval challenges, are included in the dataset completely off topic documents as well as incomplete but contextually relevant references. As for the answers, those with perfect expected marks were written from scratch, and then slightly modified to match the other test types, sometimes with the help of an AI writing assistant, but always with final human corrections. The GroUSE dataset was constructed by a single annotator who speaks fluent English.
@@ Appendix B Annotation procedures
Six annotators were asked to assess the relevancy, completeness, usefulness, and faithfulness of answers. Each annotator was tasked with evaluating all 16 answers for one or more questions, depending on their availability. If they annotated several sets, the samples of the sets annotated subsequent to the first were shuffled. They were provided with definitions of each metric adapted from GPT-4 prompts, along with the question, the references, a reference answer, and the answer to be evaluated. All annotators are fluent English speakers, one of them is familiar with the evaluated task, two have general knowledge of RAG, while the remaining three had no prior knowledge of the subject.
@@
Appendix C Prompt templates used for evaluating a grounded answer
@@
The prompts used for GPT-4 are available on Figures 7, 8, 9 and 10.
@@
Test type Test name Goal of the test Failure mode References content Answer content Answer relevancy Complete-ness Usefulness Faith-fulness Positive Accept. Negative Rejection
1 Highest marks 1 A correct answer should receive good grades. References contain a precise response. Answer contains correct response. ${}=5$ ${}=5$ ${}=\operatorname{None}$ ${}=1$ ${}=\operatorname{None}$ ${}=\operatorname{None}$
2 Highest marks 2 A correct adversarial answer with no related information should receive good grades. References contain no response but related information. Answer claims there is no response in the references. ${}=\operatorname{None}$ ${}=\operatorname{None}$ ${}=\operatorname{None}$ ${}=\operatorname{None}$ ${}=1$ ${}=1$
3 Highest marks 3 A correct adversarial answer, providing related information should receive good grades. References contain no response but related information. Answer claims there is no response in the references and completes with related information. ${}=\operatorname{None}$ ${}=\operatorname{None}$ ${}=1$ ${}=1$ ${}=1$ ${}=1$
4 Highest marks 4 A dense answer with lots of relevant information retrieved should receive good grades. References contain lots of precise information. Answer gives a correct response, more densely than the ground truth. ${}=5$ ${}=5$ ${}=\operatorname{None}$ ${}=1$ ${}=\operatorname{None}$ ${}=\operatorname{None}$
5 Highest marks 5 Same as Highest mark 2 but the references contain no related information. References are completely off-topic. Answer claims there is no response in the references. ${}=\operatorname{None}$ ${}=\operatorname{None}$ ${}=\operatorname{None}$ ${}=\operatorname{None}$ ${}=1$ ${}=1$
6 Highest marks 6 Checks that a model does not use its internal knowledge to evaluate the plausibility of the answer. The references contain absurd information which answers the question. Answer contains correct response, quoting absurd information. ${}=5$ ${}=5$ ${}=\operatorname{None}$ ${}=1$ ${}=\operatorname{None}$ ${}=\operatorname{None}$
7 Highest marks 7 Same as Highest mark 6, but for related information. The references contain no answer, but some absurd information related to the question. Answer claims there is no response in the references and completes with absurd related information. ${}=\operatorname{None}$ ${}=\operatorname{None}$ ${}=1$ ${}=1$ ${}=1$ ${}=1$
8 Low answer relevancy 1 Answer relevancy should be low when the answer contains irrelevant information. FM1 References contain a precise response. Answer contains correct response, but also irrelevant information. ${}<5$ ${}=5$ ${}=\operatorname{None}$ ${}=1$ ${}=\operatorname{None}$ ${}=\operatorname{None}$
9 Low answer relevancy 2 Answer relevancy should be minimal when the answer contains no relevant information. FM1 FM2 References are completely off-topic. Answer contains only irrelevant information, quoting the off-topic references. ${}=1$ ${}=\operatorname{None}$ ${}=\operatorname{None}$ ${}=1$ ${}=\operatorname{None}$ ${}=0$
10 Low completeness 1 Completeness should be low when the answer lacks relevant information. FM3 References contain a precise response. Answer contains most of the relevant information, but some is missing. ${}=5$ ${}<5$ ${}=\operatorname{None}$ ${}=1$ ${}=\operatorname{None}$ ${}=\operatorname{None}$
11 Low completeness 2 Completeness should be minimal when the answer wrongly claims there is no answer. FM3 FM4 References contain a precise response. Answer claims there is no response in the references. ${}=\operatorname{None}$ ${}=1$ ${}=\operatorname{None}$ ${}=\operatorname{None}$ ${}=0$ ${}=\operatorname{None}$
12 Low completeness 3 Same as above, even if the answer provides related information. FM3 FM4 References contain a precise response. Answer claims there is no response in the references and completes with related information. ${}=\operatorname{None}$ ${}=1$ ${}=1$ ${}=1$ ${}=0$ ${}=\operatorname{None}$
13 Low usefulness 1 Usefulness should be low when an answer provides unrelated information. FM5 References are completely off-topic. Answer claims there is no response in the references and completes with off-topic information. ${}=\operatorname{None}$ ${}=\operatorname{None}$ ${}=0$ ${}=1$ ${}=1$ ${}=1$
14 Low faithfulness 1 Faithfulness should be low when the answer contains an incorrect citation. FM6 References contain a precise response. Answer contains correct response, but contains a mistake in the citations. ${}=5$ ${}=5$ ${}=\operatorname{None}$ ${}=0$ ${}=\operatorname{None}$ ${}=\operatorname{None}$
15 Low faithfulness 2 Faithfulness should be low when the answer misses a citation. FM7 References contain a precise response. Answer contains correct response, but forgets one of the citations. ${}=5$ ${}=5$ ${}=\operatorname{None}$ ${}=0$ ${}=\operatorname{None}$ ${}=\operatorname{None}$
16 Low faithfulness 3 Faithfulness should be low when the response distorts the content of the references. FM7 References contain a precise response. Answer contains correct response, but distorts the content of the references. ${}=5$ ${}=5$ ${}=\operatorname{None}$ ${}=0$ ${}=\operatorname{None}$ ${}=\operatorname{None}$
Table 6: Detailed list of unit test types: their goals, which failure modes they cover, the characteristics of the references and answer, as well as the expected notes. The first seven tests cover no failure modes but check that correct answers get the highest marks possible, in various situations.
Appendix D Detailed unit test results
@@
Appendix E Dataset constitution
Fine-tuning prompt format.
@@ Appendix E Dataset constitution
Appendix F Training and inference hyperparameters
@@
Appendix G Finetuning results
@@
Appendix H Ablation: balancing the training dataset to reduce judgement biases

### v2

@@ Abstract
Sacha Muller  António Loisonthanks: For inquiries, please contact these authors.  Bilel Omrani  Gautier Viaud11footnotemark: 1 Illuin Technology { sacha.muller, antonio.loison, bilel.omrani, gautier.viaud }@illuin.tech.fr
@@ 2 Related work
Several prior works have explored methods for evaluating the generator module in RAG systems. Various sets of metrics have been proposed to measure different failure modes, but there is no consensus on a common set of criteria for evaluating grounded question answering. Chen et al. (2024) propose and evaluate 4 abilities required for RAG: noise robustness, negative rejection, information integration and counterfactual robustness. Es et al. (2023) propose two other criteria: faithfulness and answer relevancy. Faithfulness evaluates the factual consistency of the answer given the grounding contexts by decomposing the answer into several statements and calculating the proportion of facts that are supported by the contexts. Answer relevancy measures how well the provided answer addresses the original question. An LLM is used to generate several questions from the answer, the answer relevancy score is then given by the average cosine similarity between dense embeddings of the generated questions and the original question. In Deepeval [^2] https://github.com/confident-ai/deepeval, answer relevancy is computed by using a judge LLM to divide the answer into several atomic facts and computed as the proportion of facts that are relevant to the question. Yu et al. (2024) survey several prior works and propose to add correctness to this pair of metrics, which measures the accuracy of the generated response against a ground truth response. Magesh et al. (2024) focus on the legal domain and propose a more fine-grained measure of faithfulness, by distinguishing between the factual accuracy of the response and the validity of the accompanying citations. Thakur et al. (2023) introduce NoMIRACL, a human-labeled dataset of multilingual queries and both relevant and non-relevant subsets to evaluate if the generator correctly refrains from answering with non-relevant passages and correctly recognizes the relevant passage otherwise.
@@ 7. FM7
Table 1: Equivalent failure modes studied and reported in prior works. Existing studies focus on detecting and evaluating a subset of failure modes. For instance FM1 is related to answer relevancy in Es et al. (2023), FM2 is related to negative rejection in Chen et al. (2024) FM6 and FM7 are related to faithfulness and more specifically to correctness and groundedness respectively in Magesh et al. (2024), FM6 is related to citation recall in Gao et al. (2023).
@@ 4.2 Meta-evaluation with unit-testing
We introduce GroUSE (Grounded QA Unitary Scoring of Evaluators), a benchmark consisting of 144 tests divided into 9 sets of 16 tests (Figure 1). All tests within a given set share the same question, with the references and answers slightly modified to fit each of the 16 test types. An additional set of 16 tests is available as a “training set” to assist in engineering the prompt for the judge model being tested. The references are primarily excerpts from Wikipedia, while the themes of the sets span various domains, including history, science, zoology, cinematography, and the medical field. See Appendix A and B for details.
@@ 4.3 Evaluating existing Answer Relevancy and Faithfulness implementations
[…]
Figure 4: Characteristics of the 16 test types. Types 1 to 7 don’t correspond to any failure mode as they test in various situations the ability of the model to correctly evaluate answers that deserve the highest notes.
@@ 4.4 Enhancing existing frameworks
Each prompt was engineered to fit the specific metric being evaluated, but for all metrics we ask the model to rate two answers, the first one being a reference answer. The model’s reasoning is also guided by the expected format of the JSON output. Following best practices recommendations from Biderman et al. (2024), details about the prompts format and the prompt engineering process are available in Appendix D.
@@
Table 3 shows the performances of various models on GroUSE. GPT-4 is the best automatic evaluator, with an overall score of $95\%$ (very close to human performance), while the best open-source evaluator is Llama-3 70b with a score of $79\%$. The gap between closed and open-source models thus remains wide, with a $15.85$ p.p. difference on the tests results. Interestingly, Prometheus 2 8x7b does not outperform Mixtral 8x7b, despite Prometheus 2 being specialized in evaluation tasks. It’s worth noting however that Prometheus 2 was trained to predict Likert scores ranging from 1 to 5, whereas our evaluation metrics include boolean and nullable scores, which is outside its intended scope.
@@
Appendix C protocol Human annotators 98.26 92.36 97.92 95.49 96.53 96.88 96.24
@@ 5.1 Experimental setup
We finetuned a Llama-3 8b on 1k samples of this dataset, and used the rest as a test set. We refer to Appendix H for details on the finetuning configuration. To measure the model’s progress, we tested its performances on GroUSE. Following Kim et al. (2023), we also report the correlation between GPT-4’s grades and the finetuned model’s grades on the test set. For metrics using a Likert scale, alignment is measured using the Spearman correlation. Other metrics are measured using nullable boolean values, their alignment is evaluated using a three-class macro F1-score.
@@ 5.2 Experimental results
Table 3 presents the pass rate of various judge models on GroUSE, including Llama-3 8b (zero-shot) and our finetuned Llama-3 8b judge model. Finetuning significantly enhances the evaluation capabilities of Llama-3, as evidenced by the substantial improvement in pass rates. Despite extensive prompt engineering and intermediate CoT reasoning (see Appendix D), the non-finetuned Llama-3 8b passes only 40% of the unit tests. However, after finetuning, its pass rate increases to 83%, surpassing all other open-source judges, including Prometheus 2 8x7b (except in the category of faithfulness), despite its smaller size.
@@ Acknowledgments
This research work is supported by Illuin Technology. We would like to express our gratitude to Stanislas Dozias for the insightful discussions and exchange of ideas that contributed to the development of this paper. Special thanks to Quentin Lutz and Manuel Faysse for their valuable feedback and support throughout the research process. Additionally, we would like to thank Noa Grollimund, Florian Muller, Benoît Muller, Yosr Jelassi, Max Conti, and Paul Boulenger for their assistance in evaluating the human performance on GroUSE.
@@ Appendix A Unit test characteristics
The queries of the sets are:
@@
Appendix B GroUSE Dataset creation
Initial corpus creation.
We randomly selected $50$ Wikipedia pages from the 200 000200000200\,000200 000 most popular entries and scraped their content. Each page was divided into text chunks, which were subsequently clustered based on topic similarity. For each cluster, GPT-3.5 was used to generate a question that could be answered using the cluster’s content. GPT-4 was then employed to create grounded QA answers to these questions. From these generated grounded QA samples, we selected 10 examples where answering the question required synthesizing information from multiple sources rather than extracting simple facts.
Manual enhancement.
Questions were occasionally refined to encourage more complex and analytical responses. The grounding documents were also enriched by manually collecting additional relevant sources through web searches. These included excerpts from newspapers, interviews, popular science articles, and medical papers, encompassing both directly relevant and tangentially related materials. To simulate retrieval system noise, the manually collected documents were deliberately altered by sometimes truncating the last sentence. Off topic or poorly parsed documents from the automatic scraping process were kept in the grounding documents. For test types 6 and 7, GPT-4 was employed to generate documents containing intentionally absurd facts to ensure the evaluator does not rely on its internal knowledge to judge the plausibility of the information in the answer.
Answers creation and modification.
Gold-standard answers for test types 1 and 2 were manually written, using GPT-4-generated answers as initial drafts. Variations of these answers were then created with GPT-4 assistance to align with other test types. For example, a simple prompt was used to add superfluous information in the answer for test type 8. All generated content was systematically reviewed and corrected to ensure accuracy and quality.
The GroUSE dataset was constructed by a single annotator who speaks fluent English.
Appendix C Annotation procedures
@@
Seven annotators were asked to assess the relevancy, completeness, usefulness, and faithfulness of answers. Each annotator was tasked with evaluating all 16 answers for one or more questions, depending on their availability. If they annotated several sets, the samples of the sets annotated subsequent to the first were shuffled. They were provided with definitions of each metric adapted from GPT-4 prompts, along with the question, the references, a reference answer, and the answer to be evaluated. Each sample was annotated two times by different annotators and the performance was computed on the average of the pass rate of the two annotators. Table 5 shows the inter-agreement rate for each metric between the two annotators of each sample. The agreement rate is over 90%, which augments confidence about the human evaluation. All annotators are fluent English speakers, one of them is familiar with the evaluated task, three have general knowledge of RAG, while the remaining three had no prior knowledge of the subject.
@@
Inter-Agreement Rate per Metric Total Inter-Agreement rate
91.67 88.19 96.52 90.97 93.06 93.75 92.36
Table 5: Percentage of agreements between annotators when evaluating human performance.
Appendix D Prompt templates used for evaluating a grounded answer
@@
The prompts used for GPT-4 are available on Figures 8, 9, 10 and 11.
@@
Appendix E Detailed unit test results
@@
Appendix F French GroUSE Evaluation
Following a reviewer’s advice, we translated GroUSE to French to study how the performance of Judge LLMs varies when the language of the question, contexts and answers is changed, while maintaining the evaluation instructions in English. Some prompts were slightly changed to adapt to the new dataset and achieve a satisfactory score on the training set of French unit tests. As shown in Table 7, the results are generally slightly lower compared to those presented in Table 3. This performance degradation may be attributed to the predominant English data seen during the training of these models and/or the linguistic mismatch between the instruction language and the context language. Despite this variation, OpenAI models continue to demonstrate superior performance, with GPT-4-turbo emerging as the top performer in terms of total pass rate. Among open-source alternatives, Mixtral 8x22b Instruct shows promising capabilities.
Agreement rate of metrics Total test pass rate
Closed-source GPT-4 93.06 83.33 100 91.67 92.36 92.36 92.12
GPT-4o 89.58 85.42 99.31 93.75 90.97 90.97 91.67
GPT-4-turbo 90.28 88.19 99.31 89.58 95.14 95.14 92.94
GPT-3.5-turbo 85.42 50.69 80.56 66.67 70.83 64.58 69.79
Gemini 1.0 Pro 72.22 55.55 84.73 47.22 79.17 80.56 69.91
Open-source Mixtral 8x7b Instruct 77.08 52.78 83.33 68.75 78.47 74.31 72.45
Mixtral 8x22b Instruct 88.89 73.61 99.31 86.11 86.81 83.33 86.34
Llama-3.1 70b Instruct 88.19 61.11 97.92 75.00 84.03 81.94 81.37
Llama-3.1 8b Instruct 62.50 29.17 86.11 68.06 63.89 65.28 62.62
Table 7: Percentage of tests passed for various models on French samples. They were evaluated with English instruction prompts following the Figure 5 pipeline. The highest score in each column is highlighted in bold.
Appendix G Finetuning dataset constitution
Finetuning prompt format.
@@ Appendix G Finetuning dataset constitution
Appendix H Training and inference hyperparameters
@@
Appendix I Finetuning results
@@
Appendix J Ablation: balancing the training dataset to reduce judgement biases